OpenStudy (samsan9):

solve x^2+4x+5=0 and 4x^2+20x+25=0

OpenStudy (anonymous):

the equation \(\Large x^2+4x+5=0\) has no solution \(\Large 4x^2+20x+25=0 \Rightarrow x=\frac{-20\pm\sqrt{20^2-4\times4\times25}}{2\times 4}=\frac{-20}{8}=\frac{5}{2}\)